Can you please help me? My question is - I am 28 year older than my son, in 5 years time I will be double my son's age. What age am I ? Thank you.
Answer by ankor@dixie-net.com(22740)

I am 28 year older than my son, in 5 years time I will be double my son's age. What age am I ?
:
Let y = your age
let s = son's age
:
Write an equation for each statement
:
"I am 28 year older than my son,"
y = s + 28
:
"in 5 years time I will be double my son's age."
we add 5 yrs to both the ages
y + 5 = 2(s + 5)
y + 5 = 2s + 10
y = 2s + 10 - 5
y = 2s + 5
Replace y with (s+28) from the 1st statement
s + 28 = 2s + 5
28 - 5 = 2s - s
23 = s
s = 23 yrs is the son's age
and
y = 23 + 28
y = 51 yrs is your age
:
:
Check this in the 2nd statement, replace yours and sons age
"in 5 years time I will be double my son's age."
51 + 5 = 2(23 + 5)
56 = 2(28)